Problem: Translate the following into a quadratic equation, and solve it: The length of a rectangular garden is three times its width; if the area of the garden is 75 square meters, what are it's dimensions?
Answer by checkley77(12844)

area=L*W
75=3W*W
75=3W^2
W^2=75/3
W^2=25
W=SQRT25
W=5 WIDTH OF THE GARDEN.
3*5=15 FOR THE LENGTH OF THE GARDEN.
PROOF:
75=15*5
75=75